OpenStudy (anonymous):

f it takes Sue to make 4 cones for 1 sundae and and Evelyn 5 cones for 3 sundaes f sue makes 4 sundaes, how many sundaes will Evelyn make?

OpenStudy (shane_b):

If Sue makes 4 sundaes she will use 16 cones since she uses 4 cones per sundae. Evelyn uses 5 cones for every 3 sundaes and she has 16 cones to work with. Therefore the calculation would be:\[\frac{5}{3}=\frac{16}{x}\]Just solve for x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f Sue makes 4 sundaes she will use 16 cones since she uses 4 cones per sundae. Evelyn uses 5 cones for every 3 sundaes and she has 16 cones to work with. Therefore the calculation would be: 5*3=16x Just solve for x. x = 5*3/16 x = 15/16

OpenStudy (anonymous):

r8

OpenStudy (shane_b):

I'm totally confused by what you wrote. \[\frac{5}{3}=\frac{16}{x}\]\[5x=(16)(3)\]\[x=\frac{(16)(3)}{5}=\frac{48}{5}=9.6\]Since she can't make .6 cones...the can only make 9 full cones.

OpenStudy (anonymous):

ok thnks